OpenStudy (anonymous):

jim has a mass of 15 kilograms and is running on th eground with a velocity of 2.7 meters/second. What is jim's kinetic energy?

OpenStudy (anonymous):

\[E_k = \frac{ mv^2 }{ 2 }\] So mass is 15 kilograms, and the velocity is 2.7 m/s. There are already the correct SI units so no conversation necessary, simply apply the formula: \[E_k = \frac{ 15kg \times 2.7m/s }{ 2 } = \frac{ 40.5 }{ 2 } = 20.25J\]

OpenStudy (stamp):

Identify. KE = 1/2 mv^2 m = 15 kg v = 2.7 m/s Find KE — see Iamgmg90's description.

OpenStudy (anonymous):

kinetic energy=1/2mv ^{2} =1/2*15*2.7*2.7=54.675J